Subject: 3.5-m Users Committee minutes 1/14/08
From: Bruce Gillespie
Submitted: Mon, 21 Jan 2008 10:49:14 -0700
Message number: 1083
(previous: 1082,
next: 1084
up: Index)
APO 3.5-m Users Committee Phonecon, 1/14/08
Attending: Suzanne Hawley, Bruce Gillespie, Scott Anderson, Remy
Indebetouw, Al Harper, Jon Holtzman, Jon Fulbright, Mark Klaene,
Michael Strauss, Bill Ketzeback
Absent: John Bally
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - Michael said that the Princeton users were
happy. Michael was also pleased that we were able to rearrange the
schedule on short notice to accommodate special 3.5-m observations on
17 January, parallel with Chandra and VLA observing programs (the
dates for the observations were not known until about a week ago).
He's heard no feedback from NIC-FPS users regarding latent images from
bright targets.
UVa (Indebetouw) - Remy said the Virginia users are happy with their
recent observing runs.
UW (Anderson) - Scott reported that everything is fine. He's heard no
feedback from NIC-FPS users regarding latent images from bright targets.
CU (Bally) - No report.
JHU (Fulbright) - Jon had no user feedback to report.
NMSU (Holtzman) - Jon had nothing to report other than the recent
Observing Specialist training of an NMSU student seems to have gone
well. He also had not heard any feedback from users about NIC-FPS
latent images.
UChicago (Harper) - Al said that he's concerned about the recent
problems with NIC-FPS, and that his users are waiting for the NIC-FPS
Fabry-Perot to be available for science. He also mentioned that the
UChicago users have seen problems with residual images with NIC-FPS.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below. Mark highlighted the recent telescope
azimuth pointing problem, which seems to have been corrected by
remapping the drive--this is a new problem that will be watched. The
NIC-FPS ion pump was replaced. We lost fiber communications to the
upper level of the telescope this past weekend; spare fibers were put
into service, but the fiber communications system to the telescope is
in need of a wholesale overhaul due to its decrepitude. We held a CDR
on moving Agile to a new mirror-cell port, and giving it its own
rotator. The move and rotator should be completed in time for use
sometime during Q2 or Q3. The Agile focal reducer has been
reinstalled with new AR coatings, and the new throughput measurements
have been taken, which seem to indicate an improvement.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 12/06/07
through 1/08/08
Mark Klaene & Bruce Gillespie
0) Overview
Operations continued through the holiday period with minimal
disruptions to telescope, instrument, and site activities. Snowfalls
to date have been light and have not caused any significant loss of
observing time from buildup on the roof. Useful data was obtained
during engineering time that helped correct azimuth pointing errors.
1) Telescope
The azimuth pointing errors we had recently been seeing appear to have
been caused by a poor fiducial map. Once the azimuth axis was
remapped and a new pointing model obtained, pointing was improved.
There were no other issues with the telescope this period.
2) Instruments
DIS, SPIcam, and the Echelle have all operated smoothly throughout the
period, with no reported problems except one brief ion pump current
excursion on the DIS blue side. The performance of the NIC-FPS ion
pump got increasingly worse, which caused us to shut down the
instrument for approximately 6 days/nights. A new ion pump was
installed and the instrument was returned to service. The NIC-FPS
etalon and detector noise testing at the site was completed--it
appears the etalon needs to be removed for further testing, and a new
fan-out board will be manufactured to hopefully address the detector
noise issue. The newly coated focal reducer for Agile is being tested
for throughput; preliminary data reduction and analysis are underway.
The visiting instruments (CorMASS, GFP, and APOLLO) were all
operational.
3) CIF projects
A CDR was held for the implementation of Agile field rotation and its
migration to the TR2 port. Testing of the new axis controllers
continued.
**********************************
NIC-FPS status - Hawley:
Suzanne said she had recently talked with John Bally at the AAS
meeting, who told her that the etalon is not workable at present
because it cannot achieve parallelism of the plates. The plan is to
remove the etalon and return it to Boulder for further testing,
possibly next month. The new fan-out board for the detector (to
hopefully reduce detector noise) will be installed at the same time.
Al Harper thought that the problem with the present fan-out board is
that it may need more ground-plane shielding. Suzanne asked about
user feedback on residual images from bright targets. We want to make
sure that the NIC-FPS users are individually polled to see if they are
having problems with the residual images, in order to understand the
scheduling implications. The action item from last meeting will stay
open until our next meeting.
**********************************
TripleSpec and CorMASS status, plans - Indebetouw:
Remy said that TripleSpec was about to undergo its first end-to-end
test. A March delivery to APO is anticipated, and UVa would be
willing to try to support science observations on a shared-risk basis
with ARC users, assuming that the instrument is in a state to attempt
them. Suzanne suggested that we should try to collect science test
targets from the ARC users during Q2. Remy said this should be
workable, perhaps later in Q2. Suzanne noted that the Q3 proposals
need to be submitted in May, the middle of Q2, so we'll need to keep
the user community posted on the TripleSpec status as it is
commissioned. Bruce said that people who want to try shared-risk
science observing in Q2 should send their target lists to John Wilson
at UVa in the March/April timeframe. Suzanne said that she
anticipates sending an opportunity call for science targets for
TripleSpec commissioning once the instrument is delivered and we see
how it is working. UVa is also providing a data-reduction s/w package
for TripleSpec, which will also need to be exercised early. CorMASS
(which is not being used much, if at all) will stay at the site at
least through Q2 (and perhaps longer) but CorMASS will have to be
moved off the observing level when TripleSpec arrives to minimize
traffic congestion. TripleSpec will effectively replace CorMASS when
it becomes operational.
**********************************
ReSTAR report - Hawley:
Suzanne has been on the Renewing Small Telescopes for Astronomical
Research committee for the past year as an ACCORD liaison, and the
ReSTAR final report has been posted at the NOAO website (see http://www.noao.edu/)
. It outlines a national program combining federal and private
telescopes in the 2- to 4-m aperture range. The goal is to offer
approximately 8 such telescopes with oversubscription rates of ~2 or
less, and equipped with state of the art instruments for high- and low-
resolution optical spectroscopy, IR spectroscopy, and optical and IR
imaging (not every capability would be offered on each telescope
necessarily). There is an extensive science case provided in the
document. Operating these telescopes as part of a "Telescope System"
would enable NSF support. With the news that the UK is pulling back
from Gemini, there will likely be a ReSTAR-like committee formed to
discuss telescope time needs on large telescopes in the near future,
which might delay its implementation on the mid-sized telescopes. But
there is a chance that ReSTAR could be in place for the mid-sized
telescopes fairly soon. Suzanne recommended that we wait for a few
months to see what the NSF has to say about the report, and in the
meantime it would be useful for the Users Committee and interested
users to read the report.
**********************************
Mythbusters at APO [added in proof] - Hawley:
Suzanne thought it would be worthwhile to mention in the minutes that
filming for the Discovery Channel show "Mythbusters" occurred this
week at APO. TV crews shot on-site footage of APOLLO lunar ranging
while actual lasing was taking place; the show intends to discredit
the "myth" that astronauts never landed on the moon. The final
production will be aired at some future date.
**********************************
ACTION ITEMS:
[open from previous months]:
ACTION: Committee members to look at the new user webpages, and
report back to Jon Holtzman and Gabrelle Saurage suggestions for
content and organization.
STATUS: Closed. Little additional feedback. New webpages have now
become the default, with links and archive access provided to old pages.
ACTION: John Bally will write a section with figures on Fowler
sampling for inclusion in the NIC-FPS user documentation.
STATUS: Open.
ACTION: User Committee members should find out if their NIC-FPS users
are having problems with residual images caused by bright targets.
STATUS: To remain open until 2/25 meeting.
[new actions from this month]:
none
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 25 February, at
8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1083 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO